Erin's test scores were consecutive odd integers. If her next test score is 18 points more than the highest score of the three tests, her total number of points will be 328. Find all four of Erin's test scores
Answer by robertb(5830)

x + (x+2) + (x+4) + (x + 22) = 328,
4x + 28 = 328,
4x = 300,
x = 75,
x + 2 = 77,
x + 4 = 79,
x+22 = 97.